Problem #5

Admission by Investment of Assets


Geron, Aglugub, and Onate have equities in a partnership of P500,000, P800,000, and P700,000, respectively, and share profits and losses in a ratio of 5:3:2, respectively. The partners have agreed to admit Retada to the partnership.


Required: Prepare entries the journal entries to record the admission of Retada to the partnership under each of the following assumptions:


1. Retada invested P400,000 for a 25% interest, and bonus is recorded for Retada.

2. Retada invested P800,000 for a 20% interest, and bonus is recorded for the old partners.

1
Expert's answer
2021-06-01T14:01:27-0400

Solution:

1.). Existing partners capital = 500,000 + 800,000 + 700,000 = 2,000,000

New partners investment (Retada) =                                             400,000

Paid in capital after admission of Retada =                                2,400,000

New partner share = "25\%"

New partner capital = "2,400,000\times 25\% = 600,000"


Bonus to Retada calculation:

600,000 – 400,000 = 200,000

The new partner (Retada) invested 400,000 in return for a capital allocation of 600,000

Bonus to Retada = 600,000 – 400,000 = 200,000


Allocation of bonus payment to Retada from old partners:

Geron = "\frac{5}{10} \times 200,000 = 100,000"


Aglugub = "\frac{3}{10} \times 200,000 = 60,000"


Onate = "\frac{2}{10} \times 200,000 = 40,000"

The journal entry to record the admission of Retada to the partnership with bonus payment to Retada is as follows:










2.). Existing partners capital = 500,000 + 800,000 + 700,000 = 2,000,000

New partners investment (Retada) =                                              800,000

Paid in capital after admission of Retada =                                 2,800,000

New partner share = "20\%"

New partner capital = "2,400,000\times 20\% = 560,000"


Bonus to old partners calculation:

800,000 – 560,000 = 240,000

The new partner (Retada) invested 800,000 in return for a capital allocation of 560,000

Bonus to old partners = 800,000 – 560,000 = 240,000


Allocation of the bonus payment to old partners:

Geron = "\frac{5}{10} \times 240,000 = 120,000"


Aglugub = "\frac{3}{10} \times 240,000 = 72,000"


Onate = "\frac{2}{10} \times 240,000 = 48,000"


The journal entry to record the admission of Retada to the partnership with bonus payment to old partners is as follows: